OpenStudy (anonymous):

describe the difference betweem the graph of \[ y=-\left(\begin{matrix}3 \\ 4\end{matrix}\right)(x-6)^2+4\]

OpenStudy (anonymous):

and the standard position graph of y = x2

OpenStudy (anonymous):

Well tha amplitude is negative so you will have a reflection in the x axis, secondly the -6 inside the parathensis will translate your graph towards 6 units the right and the 4 will translate the graph 4 units up. The \[\left(\begin{matrix}3 \\ 4\end{matrix}\right)\] is just a scare tactic, just makes the graph really wide since \[\left(\begin{matrix}3 \\ 4\end{matrix}\right) = 3!/4!= 1/4\] Peace, Love and Happiness from Puerto Rico